# How Uniform Random Weights Induce Non-uniform Bias: Typical Interpolating Neural Networks Generalize with Narrow Teachers

### v1

@@ A.4 Relationship to PAC Bayes
$O\left(\frac{-\log\left(\tilde{p}\right)+\log\left(\frac{1}{\delta}\right)}{\epsilon}\right)$
@@
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\;\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]<\epsilon\;\right)\geq 1-\delta.$
@@
Corollary A.10.
Let $\epsilon>0$ and $\delta\in\left(0,1\right)$. Then with sample complexity
$O\left(\frac{-\log\left(\tilde{p}\right)+\log\left(\frac{1}{\delta}\right)}{\epsilon\delta}\right)$
@@
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\;\mathbb{P}_{h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)<\epsilon\right)\geq 1-\delta\;\right)\geq 1-\delta.$
@@
Using Prop. A.9,
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\;\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]<\epsilon\delta\right)\;\geq 1-\delta\,.$
@@
$\mathbb{P}_{h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\geq\epsilon\right)\leq\frac{\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]}{\epsilon}<\frac{\epsilon\delta}{\epsilon}=\delta\,.$
@@
Note that the sample complexity in Corollary A.10 is larger than the one in Corollary 3.3. Instead of additive $\log\left(\frac{1}{\delta}\right)$ factors in Corollary 3.3, here we have a multiplicative $\frac{1}{\delta}$ factor.
A.5 Proving a Refined Version of Theorem 3.2
Motivation. Note that two sources of randomness affect the sample complexity in Theorem 3.2: the random sampling of hypotheses from the prior $\mathcal{P}$ in G&C algorithm and the random sampling of the dataset from $\mathcal{D}^{N}$. To understand how each of these sources affects the obtained random complexity, we derive a refined generalization bound:
Theorem A.11 (G&C Generalization, refined).
Let $\epsilon,\delta_{\mathcal{S}}\in\left(0,1\right)$, and $\delta_{h}\in\left(0,\frac{1}{5}\right)$, and assume that $\tilde{p}<\frac{1}{2}$. For any $N$ larger than
$\left(-\log\left(\tilde{p}\right)+\log\left(\frac{1}{\delta_{\mathcal{S}}}\right)+2\log\left(\log\left(\frac{1}{\delta_{h}}\right)\right)\right)\frac{1}{\epsilon}\,,$
@@
$\mathbb{P}_{\mathcal{A}_{\mathcal{P}}}\left(\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(\mathcal{A}_{\mathcal{P}}(S))<\epsilon\right)\geq 1-\delta_{\mathcal{S}}\right)\geq 1-\delta_{h}\,.$
Remark A.12.
@@
Discussion. Theorem A.11 guarantees generalization with probability at least $1-\delta_{h}$ over the hypothesis sampling and probability $\delta_{\mathcal{S}}$ over the data sampling. This separation between $\delta_{h}$ and $\delta_{\mathcal{S}}$ highlights how both sources of randomness play a role in generalization. Interestingly, the sample complexity term $N$ exhibits a logarithmic dependence on $\delta_{\mathcal{S}}$ and only a doubly logarithmic dependence on $\delta_{h}$. Thus, for any $\delta_{\mathcal{S}}$ and $\epsilon$, the probability of not sampling a PAC interpolator decays extremely fast (doubly exponential in $N$):
$\delta_{h}=\exp\left(-\exp\left(\frac{1}{2}\left(\epsilon N+\log\left({\tilde{p}}\right)+\log\left({\delta_{\mathcal{S}}}\right)\right)\right)\right)\,.$
@@
$\mathbb{P}_{\mathcal{A}_{\mathcal{P}}}\left(\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(\mathcal{A}_{\mathcal{P}}(S))\leq\epsilon\right)\geq 1-\delta_{\mathcal{S}}\right)$
@@
$+$ $\mathbb{P}_{\mathcal{A}_{\mathcal{P}}}\left(\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(\mathcal{A}_{\mathcal{P}}(S))\leq\epsilon\right)\geq 1-\delta_{\mathcal{S}}\middle|h^{\star}\notin\mathcal{H}_{\tau}\right)\mathbb{P}_{\mathcal{A}_{\mathcal{P}}}\left(h^{\star}\notin\mathcal{H}_{\tau}\right)$
@@
This holds for any $N\geq\frac{\log\left(\frac{\tau}{\delta_{\mathcal{S}}}\right)}{\epsilon}$. Specifically, by Lemma A.7, it holds for any $N$ larger than
$\frac{\log\left(\frac{1}{\tilde{p}}\right)+\log\left(\frac{1}{\delta_{\mathcal{S}}}\right)+\log\left(\log\left(\frac{1}{\delta_{h}}\right)\right)}{\epsilon}+\frac{\tilde{p}}{\epsilon\log\left(\frac{1}{\delta_{h}}\right)}\,,$
@@
$\left(\log\left(\frac{1}{\delta_{\mathcal{S}}}\right)+2\log\left(\log\left(\frac{1}{\delta_{h}}\right)\right)-\log\left(\tilde{p}\right)\right)\frac{1}{\epsilon}\,.$
A.6 Proofs Using Nonuniform Learnability
In the next pages, we will show a result that does not use the teacher assumption for the generalization of randomly sampled networks. The result is more general than Theorem 3.2 and A.11, which were both tailored for the teacher assumption. However, the price to pay for relaxing this assumption is that the following result is slightly less tight.
Instead of the teacher assumption, we rely on the probability of interpolation, defined as:
Definition A.13.
For some random hypothesis $h$ from prior $\mathcal{P}$, the interpolation probability is defined as
$\hat{p}\triangleq\mathbb{P}_{h\sim\mathcal{P}}\left(\mathcal{L}_{\mathcal{S}}(h)=0\right)$
Theorem A.14 (Generalization of Guess & Check, restated).
Under the assumption that $\hat{p}<\frac{1}{2}$, for any $\delta,\eta\in\left(0,1\right)$, and $N\in\mathbb{N}$, we have with probability at least $1-\eta$ over $h\sim\mathcal{P}_{\mathcal{S}}$ that:
$\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\epsilon\right)\geq 1-\delta\,,$
@@ A.6 Proofs Using Nonuniform Learnability
$\epsilon=\epsilon_{\delta,\eta}=\frac{\log\left(\frac{1}{\hat{p}}\right)+\log\left(\frac{4}{\delta}\right)+\log\left(\log\left(\frac{2}{\eta}\right)\right)+2\log\log\left(\frac{\log\left(\frac{2}{\eta}\right)}{\hat{p}}+1\right)}{N}\,.$ (21)
Theorem A.14 proof sketch: We first show that
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{S}}(h_{t})=0\textrm{ and }\mathcal{L}_{\mathcal{D}}(h_{t})>\tilde{\epsilon}_{t}\right)\leq\delta_{t}\,,$
where we define
$\tilde{\epsilon}_{t}=\frac{\log{1/\delta_{t}}}{N}\,.$
@@
$\tilde{\epsilon}_{t}=\frac{\log t+2\log\log\left(t+1\right)+\log{4/\delta}}{N}\,,$
@@
$\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(\exists_{t}\left(L_{S}(h_{t})=0\textrm{and }L_{D}(h_{t})>\tilde{\epsilon}_{t}\right)\right)\leq\delta\,.$
Importantly, since this holds for any $h_{t}$ we can use our $h_{t}$ sequence from Algorithm 1 and for $t=T$ to get with probability at least $1-\delta$ over $S\sim\mathcal{D}^{N}$ that
$\mathcal{L}_{\mathcal{D}}(\mathcal{A}_{\mathcal{P}}(\mathcal{S}))\leq\tilde{\epsilon}_{T}\,.$
Finally, we use the fact that $T\mid\mathcal{S}$ is a geometric random variable with success parameter $\hat{p}<\frac{1}{2}$ to obtain
$\mathbb{P}_{h_{t}}\left(T>\frac{\log 2/\eta}{\hat{p}}\right)\leq\eta\,.$
Taking the complementary of the probability above, combined with the fact that $\tilde{\epsilon}_{T}$ is a an increasing function of $T$ concludes the theorem.
For full derivation, we proceed with some lemmas before proving Theorem A.14.
Lemma A.15.
For any $\delta>0$ and any sequence $(h_{t})_{t\in\mathbb{N}}$ of hypotheses:
$\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(\exists_{t}L_{S}(h_{t})=0\textrm{ and }L_{D}(h_{t})>\tilde{\epsilon}_{t}\right)\leq\delta$
@@
$\tilde{\epsilon}_{t}=\frac{\log t+2\log\log\left(t+1\right)+\log{4/\delta}}{N}$ (22)
@@
Set $\delta_{t}=\frac{\delta}{4t\log^{2}(t+1)}$ and we will show that $\sum_{t}\delta_{t}<\delta$. From the Integral test for convergence:
@@
where by the change of variables $u=\log\left(t\right),\mathrm{d}u=\frac{\mathrm{d}t}{t}$, we solved the integral
@@
For each $h_{t}$ separately,
$\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(L_{S}(h_{t})=0\textrm{ and}L_{D}(h_{t})>\frac{\log{1/\delta_{t}}}{N}\right)$ $=\prod_{n=1}^{N}\mathbb{P}_{\mathbf{x}\sim\mathcal{D}}\left(h_{t}\left(\mathbf{x}\right)=h_{\bm{\theta}^{\star}}\left(\mathbf{x}\right)\right)$
@@
$\leq\left(1-\frac{1/\delta_{t}}{N}\right)^{N}$
$\leq\exp\left(-N\frac{1/\delta_{t}}{N}\right)$
@@
Lemma A.16.
For any $\delta>0$, and any $h_{i}$, with probability at least $1-\delta$ over $S\sim\mathcal{D}^{N}$,
$L_{\mathcal{D}}(h)\leq\tilde{\epsilon}_{T}\,.$
where $\tilde{\epsilon}_{t}$ is defined in (22) and T is defined in Algorithm 1.
@@
Lemma A.15 applies to any sequence of hypotheses, and since $S$ is independent of the sequence $h_{t}$, we can also apply it to $h_{t}$. Since the Lemma applies to all $t$, it also applies to any random $T$, even if it depends on the sample. For the $T$ used by G&C from Algorithm 1, we always have $L_{S}(h_{T})=0$, so we get the lemma from Lemma A.15. ∎
We now wish to explain the dependence on $T$ in Lemma A.16. We will do so by the following Lemma:
Lemma A.17.
For any $\eta$, and any $S$, under the assumption that $\hat{p}<\frac{1}{2}$, we have that $\mathbb{P}_{(h_{t})}\left(T>\frac{\log 2/\eta}{\hat{p}}\right)\leq\eta$
@@
Given $S\sim\mathcal{D}^{N}$, we observe that $T$ is geometric with parameter $\hat{p}$. Using the cumulative distribution function of geometric random variables, we can bound:
$\mathbb{P}_{(h_{t})}\left(T>\frac{\log\left(2/\eta\right)}{\hat{p}}\right)=\left(1-\hat{p}\right)^{\left\lfloor\frac{\log\left(2/\eta\right)}{\hat{p}}\right\rfloor}\leq\left(1-\hat{p}\right)^{\frac{\log\left(2/\eta\right)}{\hat{p}}-1}$
@@
$\mathbb{P}_{(h_{t})}\left(T>\frac{\log\left(2/\eta\right)}{\hat{p}}\right)$ $\leq\left(1-\hat{p}\right)^{\frac{\log\left(2/\eta\right)}{\hat{p}}-1}$
$\leq\left(1-\hat{p}\right)^{\frac{\log\left(\frac{\eta}{2}\right)}{\log\left(1-\hat{p}\right)}-1}$
$=\exp\left(\left(\frac{\log\left(\frac{\eta}{2}\right)}{\log\left(1-\hat{p}\right)}-1\right)\log\left(1-\hat{p}\right)\right)$
$\leq\exp\left(\frac{\log\left(\frac{\eta}{2}\right)}{\log\left(1-\hat{p}\right)}\log\left(1-\hat{p}\right)-\log\left(1-\hat{p}\right)\right)$
$=\exp\left(\log\left(\frac{\eta}{2}\right)-\log\left(1-\hat{p}\right)\right)$
$=\frac{\frac{\eta}{2}}{1-\hat{p}}\leq\eta$
Having proved Lemma A.17 and Lemma A.16 we are ready to prove our main result.

### v2

@@ B.4 Extension to Non-Interpolators
Let $\varepsilon\in\left(0,\frac{1}{2}-\varepsilon^{\star}\right)$ and $\delta\in\left(0,1\right)$. Taking $\tau=\left\lceil\frac{\log\left(\frac{\delta}{2}\right)}{\log\left(1-\tilde{p}\right)}\right\rceil$ and $N\geq\frac{1}{2\varepsilon^{2}}\log\left(2\cdot\frac{2\tau+1}{\delta}\right)$ we get
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\leq\varepsilon^{\star}+2\varepsilon\right)\geq 1-\delta\,.$
Proof.
Recall that sampling $h\sim\mathcal{P_{S}}$ is equivalent to sampling an hypothesis with the Guess&Check algorithm. Using $h^{\star}\in\mathcal{H}_{\tau}$ to denote the event that there is a teacher equivalent hypothesis sampled within the first $\tau$ samples,
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\leq\varepsilon^{\star}+2\varepsilon\right)$ $\geq\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P_{S}}}\left(\mathcal{L_{D}}\left(h\right)\leq\varepsilon^{\star}+2\varepsilon,\mathcal{L_{S}}\left(h^{\star}\right)\leq\varepsilon^{\star}+\varepsilon,h^{\star}\in\mathcal{H}_{\tau}\right)$
$\geq 1-\frac{\delta}{2}-\left(2\tau+1\right)e^{-2\varepsilon^{2}N}$
$=1-\frac{\delta}{2}-\frac{\delta}{2}$
$=1-\delta\,.$
We can simplify the lower bound on $N$ by explicitly writing $\tau$ in terms of $\tilde{p}$ in a manner similar to Lemma B.7. Specifically, assuming $\tilde{p}<\frac{1}{2}$
$\log\left(2\tau+1\right)$ $\leq\log\left(\frac{2\log\left(\frac{\delta}{2}\right)}{\log\left(1-\tilde{p}\right)}+3\right)$
$\leq\log\left(\frac{1}{\tilde{p}+\tilde{p}^{2}}\right)+\log\left(3\tilde{p}+2\log\left(\frac{2}{\delta}\right)\right)$
$\leq\log\left(\frac{1}{\tilde{p}}\right)+\log\left(2\log\left(\frac{2}{\delta}\right)\right)+\frac{3\tilde{p}}{2\log\left(\frac{2}{\delta}\right)}$
and for $\delta<\frac{1}{5}$
$\log\left(2\tau+1\right)\leq-\log\left(\tilde{p}\right)+2\log\left(\frac{2}{\delta}\right)$
so the sample complexity is
$\frac{1}{2\varepsilon^{2}}\log\left(2\cdot\frac{2\tau+1}{\delta}\right)\leq\frac{\log\left(2\tau+1\right)+\log\left(\frac{2}{\delta}\right)}{2\varepsilon^{2}}\leq\frac{-\log\left(\tilde{p}\right)+3\log\left(\frac{2}{\delta}\right)}{2\varepsilon^{2}}$
Remark B.12.
Notice that the sample complexity is quadratically dependent on the population error, as opposed to the linear dependence in the realizable case, i.e. when the teacher is assumed to be a perfect interpolator.
B.5 Relationship to PAC Bayes
As stated in Section 3, a similar result to Lemma 3.2 could be derived with PAC-Bayes analysis (McAllester, 1999), which typically focuses on the expected population error $\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]$ of a sample from the posterior. A standard PAC-Bayes bound (Langford & Seeger, 2002; McAllester, 2003) yields the following result:
Proposition B.13.
Let $\varepsilon>0$ and $\delta\in\left(0,1\right)$. Then with sample complexity
$O\left(\frac{-\log\left(\tilde{p}\right)+\log\left(\frac{1}{\delta}\right)}{\varepsilon}\right)$
@@ B.5 Relationship to PAC Bayes
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\;\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]<\varepsilon\;\right)\geq 1-\delta.$
@@
Corollary B.14.
Let $\varepsilon>0$ and $\delta\in\left(0,1\right)$. Then with sample complexity
$O\left(\frac{-\log\left(\tilde{p}\right)+\log\left(\frac{1}{\delta}\right)}{\varepsilon\delta}\right)$
@@
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\;\mathbb{P}_{h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)<\varepsilon\right)\geq 1-\delta\;\right)\geq 1-\delta.$
@@
Using Prop. B.13,
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\;\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]<\varepsilon\delta\right)\;\geq 1-\delta\,.$
@@
$\mathbb{P}_{h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\geq\varepsilon\right)\leq\frac{\mathbb{E}_{h\sim\mathcal{P}_{\mathcal{S}}}\left[\mathcal{L}_{\mathcal{D}}(h)\right]}{\varepsilon}<\frac{\varepsilon\delta}{\varepsilon}=\delta\,.$
@@
Note that the sample complexity in Corollary B.14 is larger than the one in Corollary 3.3. Instead of additive $\log\left(\frac{1}{\delta}\right)$ factors in Corollary 3.3, here we have a multiplicative $\frac{1}{\delta}$ factor.
B.6 Proving a Refined Version of Lemma 3.2
Motivation. Note that two sources of randomness affect the sample complexity in Lemma 3.2: the random sampling of hypotheses from the prior $\mathcal{P}$ in G&C algorithm and the random sampling of the dataset from $\mathcal{D}^{N}$. To understand how each of these sources affects the obtained random complexity, we derive a refined generalization bound:
Theorem B.15 (G&C Generalization, refined).
Let $\varepsilon,\delta_{\mathcal{S}}\in\left(0,1\right)$, and $\delta_{h}\in\left(0,\frac{1}{5}\right)$, and assume that $\tilde{p}<\frac{1}{2}$. For any $N$ larger than
$\left(-\log\left(\tilde{p}\right)+\log\left(\frac{1}{\delta_{\mathcal{S}}}\right)+2\log\left(\log\left(\frac{1}{\delta_{h}}\right)\right)\right)\frac{1}{\varepsilon}\,,$
@@
$\mathbb{P}_{\mathcal{A}_{\mathcal{P}}}\left(\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(\mathcal{A}_{\mathcal{P}}(S))<\varepsilon\right)\geq 1-\delta_{\mathcal{S}}\right)\geq 1-\delta_{h}\,.$
Remark B.16.
@@
Discussion. Theorem B.15 guarantees generalization with probability at least $1-\delta_{h}$ over the hypothesis sampling and probability $\delta_{\mathcal{S}}$ over the data sampling. This separation between $\delta_{h}$ and $\delta_{\mathcal{S}}$ highlights how both sources of randomness play a role in generalization. Interestingly, the sample complexity term $N$ exhibits a logarithmic dependence on $\delta_{\mathcal{S}}$ and only a doubly logarithmic dependence on $\delta_{h}$. Thus, for any $\delta_{\mathcal{S}}$ and $\varepsilon$, the probability of not sampling a PAC interpolator decays extremely fast (doubly exponential in $N$):
$\delta_{h}=\exp\left(-\exp\left(\frac{1}{2}\left(\varepsilon N+\log\left({\tilde{p}}\right)+\log\left({\delta_{\mathcal{S}}}\right)\right)\right)\right)\,.$
@@
$\mathbb{P}_{\mathcal{A}_{\mathcal{P}}}\left(\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(\mathcal{A}_{\mathcal{P}}(S))\leq\varepsilon\right)\geq 1-\delta_{\mathcal{S}}\right)$
@@
$+$ $\mathbb{P}_{\mathcal{A}_{\mathcal{P}}}\left(\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{D}}(\mathcal{A}_{\mathcal{P}}(S))\leq\varepsilon\right)\geq 1-\delta_{\mathcal{S}}\middle|h^{\star}\notin\mathcal{H}_{\tau}\right)\mathbb{P}_{\mathcal{A}_{\mathcal{P}}}\left(h^{\star}\notin\mathcal{H}_{\tau}\right)$
@@
This holds for any $N\geq\frac{\log\left(\frac{\tau}{\delta_{\mathcal{S}}}\right)}{\varepsilon}$. Specifically, by Lemma B.7, it holds for any $N$ larger than
$\frac{\log\left(\frac{1}{\tilde{p}}\right)+\log\left(\frac{1}{\delta_{\mathcal{S}}}\right)+\log\left(\log\left(\frac{1}{\delta_{h}}\right)\right)}{\varepsilon}+\frac{\tilde{p}}{\varepsilon\log\left(\frac{1}{\delta_{h}}\right)}\,,$
@@
$\left(\log\left(\frac{1}{\delta_{\mathcal{S}}}\right)+2\log\left(\log\left(\frac{1}{\delta_{h}}\right)\right)-\log\left(\tilde{p}\right)\right)\frac{1}{\varepsilon}\,.$
B.7 Proofs Using Nonuniform Learnability
In the next pages, we will show a result that does not use the teacher assumption for the generalization of randomly sampled networks. The result is more general than Lemma 3.2 and B.15, which were both tailored for the teacher assumption. However, the price to pay for relaxing this assumption is that the following result is slightly less tight. Since we do not use the teacher assumption, with some abuse of notation we use $\mathcal{D}$ to denote the joint distribution of feature-label pairs $\left(\mathbf{x},y\right)$.
Instead of the teacher assumption, we rely on the probability of interpolation, defined as follows.
Definition B.17.
For a training set $\mathcal{S}$ and a random hypothesis $h$ from prior $\mathcal{P}$, the interpolation probability is defined as
$\hat{p}_{\mathcal{S}}\triangleq\mathbb{P}_{h\sim\mathcal{P}}\left(\mathcal{L}_{\mathcal{S}}(h)=0\right)\,.$
Theorem B.18 (Generalization of Guess & Check, restated).
Under the assumption that $\hat{p}_{\mathcal{S}}<\frac{1}{2}$ for all $\mathcal{S}$, for any $\delta,\eta\in\left(0,1\right)$, and $N\in\mathbb{N}$, we have with probability at least $1-\eta$ over $h\sim\mathcal{P}_{\mathcal{S}}$ that:
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N},h\sim\mathcal{P}_{\mathcal{S}}}\left(\mathcal{L}_{\mathcal{D}}(h)\leq\varepsilon\right)\geq 1-\delta\,,$
@@ B.7 Proofs Using Nonuniform Learnability
$\varepsilon=\varepsilon_{\delta,\eta}(\mathcal{S})=\frac{\log\left(\frac{1}{\hat{p}_{\mathcal{S}}}\right)+\log\left(\frac{4}{\delta}\right)+\log\left(\log\left(\frac{2}{\eta}\right)\right)+2\log\log\left(\frac{\log\left(\frac{2}{\eta}\right)}{\hat{p}_{\mathcal{S}}}+1\right)}{N}\,.$ (22)
Theorem B.18 proof sketch: We first show that for any sequence of hypotheses $h_{t}$
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{S}}(h_{t})=0\textrm{ and }\mathcal{L}_{\mathcal{D}}(h_{t})>\tilde{\varepsilon}_{t}\right)\leq\delta_{t}\,,$
where
$\tilde{\varepsilon}_{t}\triangleq\frac{\log{1/\delta_{t}}}{N}\,.$
@@
$\tilde{\varepsilon}_{t}=\frac{\log t+2\log\log\left(t+1\right)+\log{4/\delta}}{N}\,,$
@@
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\exists_{t}\left(\mathcal{L}_{\mathcal{S}}(h_{t})=0\textrm{ and }\mathcal{L}_{\mathcal{D}}(h_{t})>\tilde{\varepsilon}_{t}\right)\right)\leq\delta\,.$
Importantly, since this holds for any $h_{t}$ we can use our $h_{t}$ sequence from Algorithm 1 and for $t=T$ to get with probability at least $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$ that
$\mathcal{L}_{\mathcal{D}}(\mathcal{A}_{\mathcal{P}}(\mathcal{S}))\leq\tilde{\varepsilon}_{T}\,.$
Finally, we use the fact that $T\mid\mathcal{S}$ is a geometric random variable with success parameter $\hat{p}_{\mathcal{S}}<\frac{1}{2}$ to obtain
$\mathbb{P}_{h_{t}}\left(T>\frac{\log 2/\eta}{\hat{p}_{\mathcal{S}}}\right)\leq\eta\,.$
Taking the complementary of the probability above, combined with the fact that $\tilde{\varepsilon}_{T}$ is a an increasing function of $T$ concludes the theorem.
For the complete derivation, we proceed with some lemmas before proving Theorem B.18.
Lemma B.19.
For any $\delta>0$ and any sequence of hypotheses $(h_{t})_{t\in\mathbb{N}}$:
$\mathbb{P}_{\mathcal{S}\sim\mathcal{D}^{N}}\left(\exists_{t}\mathcal{L}_{\mathcal{S}}(h_{t})=0\textrm{ and }\mathcal{L}_{\mathcal{D}}(h_{t})>\tilde{\varepsilon}_{t}\right)\leq\delta$
@@
$\tilde{\varepsilon}_{t}=\frac{\log t+2\log\log\left(t+1\right)+\log{4/\delta}}{N}$ (23)
@@
Set $\delta_{t}=\frac{\delta}{4t\log^{2}(t+1)}$. We first show that $\sum_{t}\delta_{t}<\delta$. Since $\delta_{t}$ is monotonically decreasing
@@
where we used the change of variables $u=\log\left(t\right),\mathrm{d}u=\frac{\mathrm{d}t}{t}$, to solve the integral
@@
For each $h_{t}$ separately, $\mathcal{L}_{\mathcal{D}}(h_{t})>\frac{\log{1/\delta_{t}}}{N}$ is a deterministic event so
$\mathbb{P}_{S\sim\mathcal{D}^{N}}\left(\mathcal{L}_{\mathcal{S}}(h_{t})=0\textrm{ and }\mathcal{L}_{\mathcal{D}}(h_{t})>\frac{\log{1/\delta_{t}}}{N}\right)$ $=\prod_{n=1}^{N}\mathbb{P}_{\left(\mathbf{x},y\right)\sim\mathcal{D}}\left(h_{t}\left(\mathbf{x}\right)=y\right)$
@@
$\leq\left(1-\frac{\log{1/\delta_{t}}}{N}\right)^{N}$
$\leq\exp\left(-N\frac{\log{1/\delta_{t}}}{N}\right)$
@@
Lemma B.20.
For any $\delta>0$, and any realization of $\left(h_{t}\right)_{t=1}^{\infty}$, with probability at least $1-\delta$ over $\mathcal{S}\sim\mathcal{D}^{N}$,
$\mathcal{L}_{\mathcal{D}}(h_{T})\leq\tilde{\varepsilon}_{T}\,.$
where $\tilde{\varepsilon}_{t}$ is defined in (23) and T is defined in Algorithm 1.
@@
Lemma B.19 applies to any sequence of hypotheses, and since $\mathcal{S}$ is independent of the sequence $h_{t}$, we can also apply it to $h_{t}$. Since the Lemma applies to all $t$, it also applies to any random $T$, even if it depends on the sample. For the $T$ used by G&C from Algorithm 1, we always have $\mathcal{L}_{\mathcal{S}}(h_{t})=0$, so we get the result from Lemma B.19. ∎
We now wish to explain the dependence on $T$ in Lemma B.20. We will do so by the following Lemma:
Lemma B.21.
For any $\eta$, and any $\mathcal{S}$, under the assumption that $\hat{p}_{\mathcal{S}}<\frac{1}{2}$, we have that $\mathbb{P}_{(h_{t})}\left(T>\frac{\log 2/\eta}{\hat{p}_{\mathcal{S}}}\right)\leq\eta$
@@
Given $S\sim\mathcal{D}^{N}$, we observe that $T$ is geometric with parameter $\hat{p}_{\mathcal{S}}$. Using the cumulative distribution function of geometric random variables, we can bound:
$\mathbb{P}_{(h_{t})}\left(T>\frac{\log\left(2/\eta\right)}{\hat{p}_{\mathcal{S}}}\right)=\left(1-\hat{p}_{\mathcal{S}}\right)^{\left\lfloor\frac{\log\left(2/\eta\right)}{\hat{p}_{\mathcal{S}}}\right\rfloor}\leq\left(1-\hat{p}_{\mathcal{S}}\right)^{\frac{\log\left(2/\eta\right)}{\hat{p}_{\mathcal{S}}}-1}$
@@
$\mathbb{P}_{(h_{t})}\left(T>\frac{\log\left(2/\eta\right)}{\hat{p}_{\mathcal{S}}}\right)$ $\leq\left(1-\hat{p}_{\mathcal{S}}\right)^{\frac{\log\left(2/\eta\right)}{\hat{p}_{\mathcal{S}}}-1}$
$\leq\left(1-\hat{p}_{\mathcal{S}}\right)^{\frac{\log\left(\frac{\eta}{2}\right)}{\log\left(1-\hat{p}_{\mathcal{S}}\right)}-1}$
$=\exp\left(\left(\frac{\log\left(\frac{\eta}{2}\right)}{\log\left(1-\hat{p}_{\mathcal{S}}\right)}-1\right)\log\left(1-\hat{p}_{\mathcal{S}}\right)\right)$
$\leq\exp\left(\frac{\log\left(\frac{\eta}{2}\right)}{\log\left(1-\hat{p}_{\mathcal{S}}\right)}\log\left(1-\hat{p}_{\mathcal{S}}\right)-\log\left(1-\hat{p}_{\mathcal{S}}\right)\right)$
$=\exp\left(\log\left(\frac{\eta}{2}\right)-\log\left(1-\hat{p}_{\mathcal{S}}\right)\right)$
$=\frac{\frac{\eta}{2}}{1-\hat{p}_{\mathcal{S}}}\leq\eta$
Having proved Lemma B.21 and Lemma B.20 we are ready to prove our main result.